# EventHallusion: Diagnosing Event Hallucinations in Video LLMs

### v3

@@ Abstract
Recently, Multimodal Large Language Models (MLLMs) have made significant progress in the video comprehension field. Despite remarkable content reasoning and instruction following capabilities they demonstrated, the hallucination problem of these VideoLLMs is less explored compared with its counterpart in the image domain. To mitigate this gap, we propose EventHallusion, a novel benchmark that focuses on assessing the VideoLLMs’ hallucination toward event, the crux of video analysis. From a hallucination attribution perspective, our EventHallusion benchmark is curated to assess a VideoLLM’s susceptibility toward language priors and vision-language biases. On the other hand, we also propose a simple yet effective method, called Temporal Contrastive Decoding (TCD), to tackle the hallucination problems of VideoLLMs. The proposed TCD method rectifies the model’s bias toward its priors during the decoding stage by comparing the original video with a modified version, in which temporal cues are disrupted. Through comprehensive evaluation of eight open-source and two closed-source VideoLLMs on the proposed EventHallusion benchmark, we observe that the open-source models suffer significantly from hallucination problems, whereas the closed-source ones perform markedly better. By further equipping open-source VideoLLMs with the proposed TCD approach, evident performance improvements are achieved across most metrics in the EventHallusion benchmark. Our codes and benchmark data are available at https://github.com/Stevetich/EventHallusion.
@@ 1 Introduction
The field of deep learning has witnessed the extraordinary advancement of Large Language Models (LLMs) [6, 36, 10] and Multimodal Large Language Models (MLLMs) [4, 28, 5, 15, 13]. Benefiting from dramatic semantic reasoning and instruction-following capabilities, LLMs and MLLMs have become versatile assistants across multiple disciplines  [13, 41]. Despite their great success, they are still trapped in the hallucination problem, where models generate plausible but factually incorrect responses [18], potentially threatening reliability and security in applying them in real-world scenarios [33, 14]. To systematically evaluate hallucinations of MLLMs, a thread of benchmarks has been developed.  [22, 11, 29]. Specifically, POPE examines dataset distributions to investigate the hallucination phenomenon in MLLMs, focusing on frequently occurring objects and common object combinations, while MME and MMBench assess hallucinations through systematic evaluations of object existence, attributes, and relations.
@@
However, these benchmarks primarily assess hallucinations within static scenes, such as those involving objects, attributes, and inter-object relations, and therefore do not effectively address the needs of emerging VideoLLMs  [31, 23]. Two recent benchmarks [37, 12] further extend their evaluation to dynamic scenes with event-related queries. In pursuit of the rigorous evaluation, both of them curate more challenging event comprehension tasks. Specifically, HallusionBench [12] reverses all frames within a video, subsequently querying the model to determine whether the original event occurs in the altered video. VideoHallucer [37] selects long videos containing multiple events and then queries the model to evaluate the chronological order of these events. Although advancing the evaluation of event hallucination, they still suffer from the following limitations. On the one hand, the frame-reversed video in HallusionBench can deviate from the distribution of natural videos, as not all events remain meaningful when played in reverse. On the other hand, merely judging the chronological order of events given in the questions fails to capture the model’s authentic ability to comprehend the detailed content of the events [^1] A more detailed analysis of event-related VQA samples in HallusionBench [12] and VideoHallucer [37] is provided in the supplementary materials due to page limitations..
@@
When delving into VideoLLMs’ event-related hallucinations, we identify two primary causes rooted in the development processes of prevalent VideoLLMs. Firstly, building upon the foundation of LLMs, VideoLLMs inherently inherit their language priors [3, 2], which manifests as an overconfidence in their language knowledge. As illustrated in the left column of Fig.1, both Video-LLaVA and ShareGPT4Video are easily misled by the frequently occurring but misleading event “person drink coffee” asked in the question. Moreover, due to the high cost associated with scaling up high-quality video-language data, prevalent VideoLLMs typically acquire multimodal knowledge from large-scale image-language datasets. However, in these datasets, specious event-related captions often arise based solely on a single image, which inevitably introduces spurious scene-event correlations. As shown in the right column of Fig.1, both Video-LLaVA and ShareGPT4Video intuitively make the conclusion of “cats are playing” when perceiving two cats within the video, without respecting the truth.
@@
Moreover, we propose a simple yet effective method called Temporal Contrastive Decoding (TCD) to reduce the event-related hallucinations of existing VideoLLMs. Without any additional training budgets, our TCD rectifies the model’s bias to its prior during the decoding stage. Technically, for each given video, we first apply the temporal downsampling operation to create a modified counterpart. Both the original video and its modified counterpart are then input into the VideoLLMs. During the decoding stage, the logits prediction of the original video is adjusted by the logits prediction of the counterpart. The rectified logits are then used to sample the next token and continue the decoding process. Within the proposed TCD, the modified counterpart disrupts the temporal cues while preserving spatial object and scene information, making its predicted logits more susceptible to bias from the model’s event-related prior knowledge. Subtracting these amplified priors in the logits space during the decoding process helps guide the model to make more reliable judgments. As a plug-and-play approach, our TCD can consistently boost existing VideoLLM’s performances on the proposed EvenHallusion as shown in Tab.2.
@@ 3.2 Evaluation Protocol
Open-ended question answering. For automatic evaluation, we feed both generated and ground-truth answers into GPT-4o, using it as a knowledgeable evaluator to determine whether the model answers the question correctly following prior benchmarks [30]. We also use accuracy as the evaluation metric. Detailed templates of prompting GPT-4o are provided in the supplementary materials due to page limitation.
4 Temporal Contrastive Decoding
To reduce the aforementioned hallucinations in VideoLLMs, the most straightforward approach is to further enrich debiased data for finetuning. However, this approach is particularly resource-intensive for collecting such videos and annotating QA pairs. To address this, we present our Temporal Contrastive Decoding (TCD) method, which effectively alleviates hallucinations in videos without requiring additional training efforts.
Motivated by the previous work [18], our TCD method aims to reduce event hallucinations by adjusting the distribution of predicted logits at each decoding step, using a contrastive counterpart that is more aligned with the model’s inherent biases. For clarity, we first review the standard auto-regressive decoding process of a VideoLLM. Given the text instruction $x$ and the video $V$, the predicted logit $z_{t}^{ori}$ at the $t$-th timestep can be formulated as:
$p_{t}=\operatorname{softmax}\Big{[}(1+\alpha)z_{t}^{ori}-\alpha z_{t}^{con}\Big{]}$ (3)
where $\alpha$ is the hyperparamter to control the contrastive strength during the decoding. However, applying such a contrastive process throughout the entire decoding stage would lead to significant model degradation, as it is not always consistent for $z_{t}^{ori}$ and $z_{t}^{con}$ to have mismatched distribution. When $z_{t}^{ori}$ and $z_{t}^{con}$ are consistent and correct, employing Eq.3 to modulate the probability impairs the model’s judgment. Therefore, we introduce another hyperparameter $\beta$ to constrain the token set used for TCD. The motivation behind this design is twofold: when the model exhibits high confidence in a specific token, we conduct TCD on a smaller subset of tokens with similarly high confidence. Conversely, when the model’s confidence is distributed more evenly across tokens, we apply a lower threshold, enabling TCD on a broader range of tokens.
@@ 4 Temporal Contrastive Decoding
Table 2: Results of TCD on EventHallusion. Default denotes the standard decoding strategy, while TCD represents our Temporal Contrastive Decoding method. We exhibit the accuracy of the binary classification task. The best results of each category are bolded.
@@ 5.1 Evaluation on EventHallusion
5.1.1 Selected VideoLLMs for Evaluation
We test 8 open-sourced Video and 2 close-sourced VideoLLMs on our EventHallusion. Open-sourced VideoLLMs comprise VideoChatGPT [32], VideoChat2 [21], PLLaVA [39], LLaMA-VID [23], VideoLaVIT [16], LLaVA-NeXT-Video [19], VILA [26], ShareGPT4Video [8], and Video-LLaVA [25]. For close-sourced VideoLLMs, we select the three most popular video-supported MLLMs, namely GPT-4V [1], GPT-4o and Gemini-1.5-Pro [35], to conduct the systematic comparison of the robustness between open-sourced and close-sourced MLLMs.
@@ 5.1.1 Selected VideoLLMs for Evaluation
5.1.2 Performance Analysis
We exhibit the performance of each VideoLLM in Tab 1 on the binary classification and detailed description matching task. As the results indicate, open-source VideoLLMs exhibit a consistent weakness in event hallucinations. In contrast, closed-source VideoLLMs demonstrate greater robustness against event hallucinations across all categories. A deeper analysis is provided below: (1) In evaluating the susceptibility of VideoLLMs to vision-language correlation biases (i.e, “Entire” and “Mix” in Tab.1), most VideoLLMs perform worse when confronted with videos containing a mix of common and rare events, compared to those containing only rare events. This result indicates that frequent events can distract the model’s attention, imposing higher challenges to the model. (2) In evaluating the susceptibility of VideoLLMs to language priors (i.e., “Misleading” in Tab.1), closed-source models demonstrate overwhelming advantages compared to open-source models, suggesting that these closed-source VideoLLMs are built upon language foundation models that are significantly more powerful than their open-source counterparts.
5.2 TCD Evaluation
5.2.1 VideoLLM Baselines.
5.2.2 Implementation Details.
5.2.3 Main Results Analysis
We exhibit the results of the binary classification task on EventHallusion in Tab.2. As shown in Tab.2, our TCD demonstrates significant performance across all three VideoLLMs, with increases in overall accuracy, confirming the effectiveness of TCD in mitigating event hallucinations on VideoLLMs with different architectures and visual-language pertaining paradigms. However, while TCD shows consistent and notable improvements on Mix and Misleading across all three models, the improvement on Entire is minimal, with even a decrease situation observed. This result suggests that entire rare events are a more challenging and hallucination-prone scenario for VideoLLMs.
5.2.4 Ablations on Contrastive Sample Constructions.
We also discuss some different alternative strategies when constructing the counterpart of the original video for performing our TCD method as shown in Tab.3. Specifically, we compare our base TCD method with adding Gaussian noises to video frames when employing TCD method (denoted as TCD+noises). As results have shown in Tab.3, While both TCD and TCD+noises demonstrate improvement on regular decoding. However, TCD+noises does not exhibit large improvements on simply using TCD method.
@@ 5.2.3 Main Results Analysis
Table 4: Ablation study of $\alpha$ in our TCD. The best results of each category are bolded.
$\alpha$ Entire Mix Misleading Overall
0.25 50.00 73.56 73.53 66.99
0.5 50.88 73.56 74.51 67.48
0.75 50.88 71.50 74.51 66.50
1.0 49.12 64.25 74.51 62.59
Table 5: Ablation study of $\beta$ in our TCD. The best results of each category are bolded.
$\beta$ Entire Mix Misleading Overall
0.1 50.88 73.58 74.51 67.48
0.2 50.88 73.58 74.51 67.48
0.5 50.88 74.09 74.51 67.73
0.5 50.00 74.61 74.51 67.73
5.2.5 Ablations on Hyper-parameters of TCD
We conduct ablation studies to investigate the effects of hyper-parameters $\alpha$ and $\beta$ on our TCD method as shown in Tab.4 and Tab.5 For the first hyper-parameter $\alpha$, with the growth of $\alpha$, TCD applies a more fierce contrast between normal and distorted video input. As results are shown in Tab.4, when $\alpha$ is 0.5, TCD exhibits best performance. While for $\beta$, the threshold used to control the token set for TCD does not exhibit a strong impact on the performance of TCD on our EventHallusion Benchmark.
5.2.6 Effect of Sampled Frames in Distorted Video.
We further explore the effect of sampled frames in distorted video on our TCD. Specifically, we adopt LLaVA-NeXT-Video as the backbone model with 32 frames sampled for the original video input and adjust the numbers of sampled frames in the distorted video. We conduct this experiment on the binary classification task of EventHallusion. As results have shown in Tab.6, TCD maintains a steady improvement in the range of 4 to 16 sampled frames. However, when the number of sampled frames in distorted video decreases from 4 to 2, the performance of TCD deteriorates. We attribute this phenomenon to the excessive destruction of the temporal modeling in the VideoLLM with too few frames incorporated. This leads to drastic changes in the logit of distorted video, thereby affecting the contrastive decoding performance.
Table 6: Ablation study of extracted video frames in the distorted group. The best results of each category are bolded.
@@ 5.3 Qualitative Results
We exhibit an example of our question in EventHallusion. As shown in Fig.6, VideoLLM can be affected and generate hallucination answers, while our TCD method can alleviate this problem. Due to the page limit, we provide more examples in our suuplementary materials.

### v4

@@ Abstract
Recently, Multimodal Large Language Models (MLLMs) have made significant progress in the video comprehension field. Despite remarkable content reasoning and instruction following capabilities they demonstrated, the hallucination problem of these VideoLLMs is less explored compared with its counterpart in the image domain. To mitigate this gap, we propose EventHallusion, a novel benchmark that focuses on assessing the VideoLLMs’ hallucination toward event, the crux of video analysis. From a hallucination attribution perspective, our EventHallusion benchmark is curated to assess a VideoLLM’s susceptibility toward language priors and vision-language biases. On the other hand, we also propose a simple yet effective method, called Temporal Contrastive Decoding (TCD), to tackle the hallucination problems of VideoLLMs. The proposed TCD method rectifies the model’s bias toward its priors during the decoding stage by comparing the original video with a modified version, in which temporal cues are ambiguited. Through comprehensive evaluation of eight open-source and three closed-source VideoLLMs on the proposed EventHallusion benchmark, we observe that the open-source models suffer significantly from hallucination problems, whereas the closed-source ones perform markedly better. By further equipping open-source VideoLLMs with the proposed TCD approach, evident performance improvements are achieved across most metrics in the EventHallusion benchmark.
@@ 1 Introduction
The field of deep learning has witnessed the extraordinary advancement of Large Language Models (LLMs) [6, 39, 10] and Multimodal Large Language Models (MLLMs) [4, 31, 5, 17, 15]. Benefiting from dramatic semantic reasoning and instruction-following capabilities, LLMs and MLLMs have become versatile assistants across multiple disciplines  [15, 44]. Despite their great success, they are still trapped in the hallucination problem, where models generate plausible but factually incorrect responses [21], potentially threatening reliability and security in applying them in real-world scenarios [36, 16]. To systematically evaluate hallucinations of ImageLLMs, a thread of benchmarks has been developed [25, 12, 32]. Specifically, POPE [25] investigates the hallucination by mining the frequently occurring objects and their combinations. MME [12] and MMBench [32] evaluate hallucinations of the existence, attributes, and relations of objects.
@@
In fact, based on the foundation of ImageLLMs, VideoLLMs have shown significant proficiency in understanding the content of single-frame images [40]. However, their effectiveness diminishes when tasked with comprehending the event, namely the actions performed by humans or objects spanning the entire video [11, 46, 19]. This limitation is further proved by our results presented in Fig.6. To further evaluate the event-oriented hallucinations of VideoLLMs, two recent benchmarks [40, 13] develop distinct event comprehension tasks. Specifically, HallusionBench [13] reverses all frames within a video, subsequently querying the model to determine whether the original event occurs in the altered video. VideoHallucer [40] selects long videos containing multiple events and then queries the model to evaluate their chronological orders. Although advancing the evaluation of event hallucination, they still suffer from the following limitations. On the one hand, the frame-reversed video in HallusionBench can deviate from the distribution of natural videos, as not all events remain meaningful when played in reverse. On the other hand, merely judging the chronological order of events given in the questions fails to probe the model’s authentic ability to comprehend the complex events [^1] A more detailed analysis of event-related VQA samples in HallusionBench [13] and VideoHallucer [40] is provided in the supplementary materials due to page limitations..
@@
When delving into VideoLLMs’ event-related hallucinations, we identify two primary causes rooted in the development processes of prevalent VideoLLMs. Firstly, building upon the foundation of LLMs, VideoLLMs inherently inherit their language priors [3, 2], which manifests as an overconfidence in their language knowledge. As shown in the left column of Fig.1, both Video-LLaVA[28] and ShareGPT4Video[8] are easily misled by the frequently occurring but misleading event “person drink coffee” asked in the question. Moreover, due to the high cost associated with scaling up high-quality video-language data, prevalent VideoLLMs typically acquire multimodal knowledge from large-scale image-language datasets. However, in these datasets, specious event-related captions often arise based solely on a single image, which inevitably introduces spurious scene-event correlations. As shown in the right column of Fig.1, both Video-LLaVA[28] and ShareGPT4Video[8] intuitively make the conclusion of “cats are playing” when perceiving two cats within the video, without respecting the truth.
@@
Additionally, we propose a simple yet effective method called Temporal Contrastive Decoding (TCD) to reduce the event-related hallucinations of existing VideoLLMs in a training-free manner. The core motivation of our TCD approach is to first elicit the spurious priors and then eliminate their adverse effects before making the final decisions. Technically, for each given video, we first construct its counterpart by ambiguiting the temporal cues while preserving the spatial information. This counterpart disentangles the event-unrelated elements that could potentially trigger spurious prior assumptions from the original video. Therefore, we feed both the original video and its counterpart into a VideoLLM and adaptively adjust the prediction distribution of the former using the latter, leading to more reliable final judgments. As a plug-and-play approach, our TCD can consistently boost existing VideoLLM’s performances on the proposed EvenHallusion as shown in Tab.2.
@@ 3.1 Data Curation
For the ground-truth answers to all types of questions, we meticulously verify them to ensure their correctness.
@@ 3.2 Evaluation Protocol
Open-ended question answering. For automatic evaluation, we feed both generated and ground-truth answers into GPT-4o[14], using it as a knowledgeable evaluator to determine whether the model answers the question correctly following prior benchmarks [33]. We also use accuracy as the evaluation metric. Detailed templates of prompting GPT-4o are provided in the supplementary materials due to page limitation.
@@
4 Temporal Contrastive Decoding
To mitigate the aforementioned hallucinations in VideoLLMs, the most straightforward solution is to enhance finetuning with additional debiased data. However, this approach is highly resource-intensive, as it requires extensive efforts to collect such videos and annotate corresponding QA pairs. Therefore, we take inspiration from [21] and propose our Temporal Contrastive Decoding (TCD) method, which alleviates event-oriented hallucinations in VideoLLMs by comparing the logits of the original video and its counterpart during the decoding stage.
For clarity, we first review the standard auto-regressive decoding process of a VideoLLM. Given the text instruction $x$ and the video $V$, the predicted logit $z_{t}^{ori}$ at the $t$-th timestep can be formulated as:
$z_{t}=(1+\alpha)z_{t}^{ori}-\alpha z_{t}^{con}$ (3)
Figure 6: Performance comparison of prevalent Video LLMs and human evaluations on object-related and event-related content recognition accuracy (i.e., our official EventHallusion evaluation) in our EventHallusion benchmark.
where $\alpha$ is the hyperparamter to control the contrastive strength during the decoding. Since $z_{t}^{ori}$ and $z_{t}^{con}$ are not always exhibited contrastively, we eliminate the adverse effects of suppressing true positives by directly filtering implausible answers with a dynamic confidence threshold $t$:
$t=\beta*\mathrm{max}(z^{ori}_{t})$ (4)
The ultimate modulated probability $p_{t}$ can be formulated as:
$p_{t}[k]=\begin{cases}\operatorname{softmax}\Big{(}z_{t}[k]\Big{)}&\text{if }z_{t}[k]>=t,\\ -\infty&\text{if }z_{t}[k]<0.\end{cases}$ (5)
where $z_{t}[k]$ denotes the $k$-th dimension of $z_{t}$.
Table 2: Results of TCD on EventHallusion. Default denotes the standard decoding strategy, while TCD represents our Temporal Contrastive Decoding method. We exhibit the accuracy of the binary classification task. The best results of each category are bolded. “LNV” is short for “LLaVA-Next-Video”.
@@ 5.1 Evaluation on EventHallusion
5.1.1 Performances of Human Evaluation
Firstly, to evaluate the quality of our proposed EventHallusion benchmark, we ask three humans to answer the questions and average their resulting scores as the human evaluation results. Meanwhile, to validate the value of evaluating event-oriented hallucinations for VideoLLMs, we also extracted object contents from videos in our EventHallusion benchmark and tested them using both human evaluations and two prevalent VideoLLMs. The above results are shown in Fig.6, where we have the following conclusions: (1) Although prevalent VideoLLMs perform poorly in EventHallusion (i.e, “Event” part in Fig.6), humans can correctly answer almost all questions. This demonstrates that the questions in our EventHallusion benchmark are clear enough to allow for accurate judgments. (2) When recognizing objects in videos, prevalent VideoLLMs can achieve human-level accuracies. However, they significantly lag behind humans when comprehending event-related content. This demonstrates that event-oriented hallucinations confuse VideoLLMs much more severely than object-related ones.
5.1.2 Peformances of Prevalent VideoLLMs
Evaluation Setup. We test 8 open-sourced Video and 3 close-sourced VideoLLMs on our EventHallusion. Open-sourced VideoLLMs comprise VideoChatGPT [35], VideoChat2 [24], PLLaVA [42], LLaMA-VID [26], VideoLaVIT [18], LLaVA-NeXT-Video [22], VILA [29], ShareGPT4Video [8], and Video-LLaVA [28]. For close-sourced VideoLLMs, we select the three most popular video-supported MLLMs, namely GPT-4V [1], GPT-4o[14] and Gemini-1.5-Pro [38], to conduct the systematic comparison of the robustness between open-sourced and close-sourced MLLMs.
@@ 5.1.2 Peformances of Prevalent VideoLLMs
Performance Analysis. We exhibit the performance of each VideoLLM in Tab 1 on the binary classification and detailed description matching task. As the results indicate, open-source VideoLLMs exhibit a consistent weakness in event hallucinations. In contrast, closed-source VideoLLMs demonstrate greater robustness against event hallucinations across all categories. A deeper analysis is provided below: (1) In evaluating the susceptibility of VideoLLMs to vision-language correlation biases (i.e, “Entire” and “Mix” in Tab.1), most VideoLLMs perform worse when confronted with videos containing a mix of common and rare events, compared to those containing only rare events. This result indicates that frequent events can distract the model’s attention, imposing higher challenges to the model. (2) In evaluating the susceptibility of VideoLLMs to language priors (i.e., “Misleading” in Tab.1), closed-source models demonstrate overwhelming advantages compared to open-source models, suggesting that these closed-source VideoLLMs are built upon language foundation models that are significantly more powerful than their open-source counterparts.
5.2 TCD Evaluation
5.2.1 VideoLLM Baselines.
Figure 7: Effects of $\alpha$ and $\beta$ in our TCD.
Table 4: Ablation study of extracted video frames when constructing the contrastive video.
@@ 5.2 TCD Evaluation
5.2.2 Implementation Details.
5.2.3 Effectiveness of TCD
We apply our proposed TCD method on three prevalent VideoLLMs, namely LLaVA-NeXT [22], VideoChat2 [35] and VILA [29], and evaluate their performances on EventHallusion. As shown in Tab.2, equipping with our TCD, three VideoLLMs achieve consistent performance boost. This demonstrates that our TCD method can be seamlessly integrated with versatile VideoLLMs and reduce their event-related hallucinations without bells and whistles.
5.2.4 Explorations on Contrastive Sample Constructions
We further explore more alternative strategies to construct the contrastive counterpart in our TCD method. Recalling that our TCD aims at ambiguating the temporal cues of the original video, we further explore the effects of blurring the spatial cues by adding gaussian noises to video frames when constructing the contrastive counterpart. As shown in Tab.3, introducing Gaussian noises on top of TCD can further boost the overall performances. However, the extent of introducing Gaussian noises should be delibrately adjusted to prevent degradation in practice. Therefore, we do not choose to apply this strategy in our TCD method.
5.2.5 Ablations on Hyper-parameters of TCD
We also ablate the effects of $\alpha$ and $\beta$ in our TCD method. As shown in Fig.7, we have the following observations: (1) With $\alpha$ increases from 0.25 to 0.5, TCD enjoys the benefits of the contrastive video’s modulation. However, further increasing $\alpha$ to 1.0 causes TCD to degrade significantly, suggesting that excessively amplifying the influence of original samples will cause the model to decode primarily according to the original distribution, rendering TCD ineffective. (2) Increasing $\beta$ from 0.2 to 0.5 enhances the model’s performance, while further decreasing or increasing $\beta$ does not result in performance fluctuations. This demonstrates that our TCD is relatively robust to the variation in $\beta$ values.
@@ 5.2.5 Ablations on Hyper-parameters of TCD
5.2.6 Effect of Sampled Frames in Distorted Video.
We further explore the effects of the temporal downsampling ratio when constructing the contrastive video in our TCD. Specifically, we adopt LLaVA-NeXT-Video[22] as the backbone model with 32 frames sampled for the original video input and downsample the number of sampled frames to construct the contrastive video. This experiment is conducted on the binary classification task of EventHallusion. From Tab.4, we have the following observations: (1) Extreme temporal downsampling (i.e, 1 frame) can lead to severe performance degradation of our TCD approach. This phenomenon may arise because single-frame input significantly diminishes spatial cues, leading to distorted prediction logits during contrastive learning. (2) A low temporal downsampling rate (e.g., 16 frames) performs worse than higher rates (e.g., 4 and 8 frames). This indicates that narrowing the gap between the contrastive video and the original one weakens the advantages of our TCD approach.
@@ 5.3 Qualitative Results
We also present qualitative results comparing the baseline method with and without our TCD approach on the EventHallusion benchmark. As shown in Fig.8, The baseline method fails to generate accurate responses, whereas incorporating our TCD approach yields correct answers. Due to the page limitation, we provide more examples in our supplementary materials.